# Zero-shot Logical Query Reasoning on any Knowledge Graph

### 10 Apr 2024 v1

@@ 4.1 Inductive Relation Projection
The vast majority of CLQA methods are inherently transductive and implement relation projections as functions over entity and relation embeddings fixed to a certain KG vocabulary, e.g., with scoring functions from KG completion methods (Hamilton et al., 2018; Arakelyan et al., 2021; Bai et al., 2023), geometric functions (Ren et al., 2020; Zhang et al., 2021b), or pure neural methods (Amayuelas et al., 2022; Wang et al., 2023). The only method inductive to new entities (Zhu et al., 2022) learns relation embeddings and uses those as a labeling trick (Section 3) for a GNN that implements the projection operator.
As fixed relation embeddings do not transfer to new KGs with new relations, in UltraQuery, we opt for Ultra (Galkin et al., 2023), an inductive approach that builds relation representations dynamically using the invariance of relation interactions. Ultra leverages theoretical findings in multi-relational link prediction (Barcelo et al., 2022; Huang et al., 2023) and learns relation representations from a meta-graph of relation interactions [^2] The meta-graph can be efficiently obtained from any KG.. The meta-graph includes four learnable edge types or meta-relations (head-to-tail, tail-to-head, head-to-head, tail-to-tail) which are independent from KG’s relation vocabulary and therefore transfer across any graph. Practically, given a graph $\mathcal{G}$} and projection query $(h,r,?)$, Ultra employs labeling trick GNNs on two levels. First, it builds a meta-graph ${\mathcal{G}}_{r}$ of relation interactions (a graph of relations where each node is a unique edge type in $\mathcal{G}$}) and applies a labeling trick to initialize the query node $r$. Running a message passing GNN over ${\mathcal{G}}_{r}$ results in conditional relation representation which are used as initial edge type features in the second, entity-level GNN. There, a starting node $h$ is initialized with a query vector from the obtained relation representations and running another GNN over the entity graph (with a final sigmoid readout) returns a scalar score in $[0,1]$ representing a probability of each node to be a tail of a query $(h,r,?)$.
Figure 2(a) illustrates the intersection-projection query execution process where each projection step is tackled by the same inductive projection operator with initialization depending on the start anchor node or intermediate variables.
@@ 4.2 Inductive Logical Operations
Learnable logical operators parameterized by neural nets in many CLQA approaches (Hamilton et al., 2018; Ren et al., 2020; Zhang et al., 2021b; Amayuelas et al., 2022) fit a particular embedding space and are not transferable. Instead, we resort to differentiable but non-parametric fuzzy logics (van Krieken et al., 2022) that implement logical operators as algebraic operations (t-norms for conjunction and t-conorms for disjunction) in a bounded space $[0,1]$ and are used in several neuro-symbolic CLQA approaches (Arakelyan et al., 2021; Zhu et al., 2022; Arakelyan et al., 2023; Bai et al., 2023; Yin et al., 2023b). UltraQuery employs fuzzy logical operators over fuzzy sets ${\bm{x}}\in[0,1]^{|{\mathcal{V}}|}$ as the relation projection operator assigns a scalar in range $[0,1]$ for each entity in a graph. The choice of a fuzzy logic is often a hyperparameter although van Krieken et al. (2022) show that the product logic is the most stable. In product logic, given two fuzzy sets ${\bm{x}},{\bm{y}}$, conjunction is element-wise multiplication ${\bm{x}}\odot{\bm{y}}$ and disjunction is ${\bm{x}}+{\bm{y}}-{\bm{x}}\odot{\bm{y}}$. Negation is often implemented as $\mathbf{1}-{\bm{x}}$ where $\mathbf{1}$ is the universe vector of all ones. For second- and later $i$-th hop projections, we obtain initial node states ${\bm{h}}_{v}$ by weighting a query vector ${\bm{r}}_{i}$ with their probability score $x_{v}$ from the fuzzy set of a previous step: ${\bm{h}}_{v}=x_{v}{\bm{r}}_{i}$.
@@ 4.3 Training
Following existing works (Ren & Leskovec, 2020; Zhu et al., 2022), UltraQuery is trained on complex queries to minimize the binary cross entropy loss.
${{\mathcal{L}}=-\frac{1}{|{\mathcal{A}}_{q}|}\sum_{a\in{\mathcal{A}}_{q}}\log p(a|q)-\frac{1}{|{\mathcal{V}}\backslash{\mathcal{A}}_{q}|}\sum_{a^{\prime}\in{\mathcal{V}}\backslash{\mathcal{A}}_{q}}\log(1-p(a^{\prime}|q))}$
@@ 5 Experiments
Our experiments focus on answering the following research questions: (1) How does a single UltraQuery model perform in the zero-shot inference mode on unseen graphs and queries compared to the baselines? (2) Does UltraQuery retain the quality metrics like faithfullness and identify the easy answers reachable by traversal? (3) How does the multi-source propagation issue affect the final performance?
@@ 5.1 Setup and Datasets
Transductive (3 datasets) where training and inference graphs are the same $(\mathcal{G}_{\textit{train}}=\mathcal{G}_{\textit{inf}})$ and test queries cover the same set of entities and relations: FB15k237, NELL995 and FB15k all from Ren & Leskovec (2020) with at most 100 answers per query.
Inductive entity $(e)$ (9 datasets) from Galkin et al. (2022b) where inference graphs extend training graphs $(\mathcal{G}_{\textit{train}}\subset\mathcal{G}_{\textit{inf}})$ being up to 550% larger in the number of entities. The set of relations is fixed in each training graph and does not change at inference making the setup inductive with respect to the entities. Training queries might have more true answers in the extended inference graph.
Inductive entity and relation $(e,r)$ (11 datasets): we sampled a novel suite of WikiTopics-QA datasets due to the absence of standard benchmarks evaluating the hardest inductive setup where inference graphs have both new entities and relations $(\mathcal{G}_{\textit{train}}\neq\mathcal{G}_{\textit{inf}})$. The source graphs were adopted from the WikiTopics datasets (Gao et al., 2023), we follow the BetaE setting when sampling 14 query types with at most 100 answers. More details on the dataset creation procedure are in Appendix A.
Implementation and Training. UltraQuery was trained on one FB15k237 dataset with complex queries for 10,000 steps with batch size of 32 on 4 RTX 3090 GPUs for 2 hours (8 GPU-hours in total). We initialize the model weights with an available checkpoint of Ultra reported in Galkin et al. (2023). Following the standard setup in the literature, we train the model on 10 query types and evaluate on all 14 patterns. We employ product t-norm and t-conorm as non-parametric fuzzy logic operators to implement conjunction $(\wedge)$ and disjunction $(\lor)$, respectively, and use a simple 1x1-x negation. For the ablation study, UltraQuery LP uses the same frozen checkpoint (pre-trained on simple 1p link prediction) with scores thresholding to alleviate the multi-source propagation issue (Section 4.1). More details on all hyperparameters are available in Appendix B.
Figure 3: Zero-shot query answering performance (MRR, higher is better) of a single UltraQuery model trained on one FB15k237 queries dataset compared to the best available baselines on 23 datasets – trainable for each transductive and inductive $(e)$ dataset, and the heuristic baseline for newly introduced inductive $(e,r)$ datasets. EPFO is the average of 9 query types with $(\wedge,\lor)$ operators, Negation is the average of 5 query types with the negation operator $(\neg)$. UltraQuery LP is an ablated version trained only on 1p link prediction. On average, UltraQuery outperforms the best baselines. More results in Table 1 and Appendix C.
@@
As common in the literature (Ren & Leskovec, 2020; Ren et al., 2023), the answer set of each query is split into easy and hard answers. Easy answers are reachable by graph traversal and do not require inferring missing links whereas hard answers are those that involve at least one edge to be predicted at inference. In the rank-based evaluation, we only consider ranks of hard answers and filter out easy ones and report filtered Mean Reciprocal Rank (MRR) and Hits@10 as main performance metrics.
Other qualitative metrics include: (1) faithfullness (Sun et al., 2020), i.e., the ability to recover easy answers reachable by graph traversal. Here, we follow the setup in Galkin et al. (2022b) and measure the performance of training queries on larger inference graphs where the same queries might have new true answers; (2) the ROC AUC score to estimate whether a model ranks easy answers higher than hard answers – we compute ROC AUC over unfiltered scores of easy answers as positive labels and hard answers as negative. (3) Mean Absolute Percentage Error (MAPE) (Zhu et al., 2022) between the number of answers extracted from model’s predictions and the number of ground truth answers (easy and hard combined) to estimate whether CLQA models can predict the cardinality of the answer set.
Baselines. In transductive and inductive $(e)$ datasets, we compare a single UltraQuery model with the best reported models trained end-to-end on each graph (denoted as Best baseline in the experiments): QTO (Bai et al., 2023) for 3 transductive datasets (FB15k237, FB15k, and NELL995) and GNN-QE (Galkin et al., 2022b) for 9 inductive $(e)$ datasets. While a single UltraQuery model has 177k parameters, the baselines are several orders of magnitude larger with a parameters count depending on the number of entities and relations, e.g., a QTO model on FB15k237 has 30M parameters due to having 2000dd entity and relation embeddings, and GNN-QE on a reference FB 175% inductive $(e)$ dataset has 2M parameters. For a newly sampled suite of 11 inductive $(e,r)$ datasets, we compare against the edge-type heuristic baseline introduced in Galkin et al. (2022b). The heuristic selects the candidate nodes with the same incoming relation as the last hop of the query. More details on the baselines are reported in Appendix B
@@
Model Inductive $(e,r)$ (11 datasets) Inductive $(e)$ (9 datasets) Transductive (3 datasets) Total Average (23 datasets) EPFO avg neg avg EPFO avg neg avg EPFO avg neg avg EPFO avg neg avg MRR H@10 MRR H@10 MRR H@10 MRR H@10 MRR H@10 MRR H@10 MRR H@10 MRR H@10 Best baseline 0.014 0.029 0.004 0.007 0.328 0.469 0.176 0.297 0.468 0.603 0.259 0.409 0.196 0.276 0.105 0.173 UltraQuery 0-shot 0.275 0.375 0.153 0.242 0.321 0.479 0.156 0.291 0.335 0.467 0.132 0.260 0.301 0.428 0.152 0.264 UltraQuery LP 0-shot 0.236 0.358 0.084 0.148 0.285 0.455 0.094 0.190 0.334 0.487 0.164 0.270 0.268 0.413 0.098 0.181 UltraQuery LP no thrs. 0.213 0.317 0.069 0.121 0.255 0.405 0.083 0.169 0.296 0.433 0.143 0.237 0.240 0.367 0.085 0.155
Figure 4: Mitigation of the multi-source message passing issue (Section 4) with UltraQuery: while UltraQuery LP (pre-trained only on 1p link prediction) does reach higher 1p query performance (center right), it underperforms on negation queries (center left). UltraQuery adapts to the multi-source message passing scheme and trades a fraction of 1p query performance for much better averaged EPFO, e.g., on the 3i query (right), and negation queries performance. More results are in Appendix C.
5.2 Main Experiment: Zero-shot Query Answering
@@
Averaged across 23 datasets, UltraQuery outperforms available baselines by relative 50% in terms of MRR and Hits@10 on both EPFO and negation queries (e.g., 0.30 vs 0.20 MRR on EPFO queries and 0.147 vs 0.105 on negation queries). The largest gains are achieved on the hardest inductive $(e,r)$ datasets where the heuristic baseline is not able to cope with the task. On inductive $(e)$ datasets, UltraQuery LP outperforms the trainable SOTA GNN-QE model on larger inductive inference graphs and performs competitively on smaller inductive versions. On transductive benchmarks, UltraQuery lags behind the SOTA QTO model which is expected and can be attributed to the sheer model size difference (177k of UltraQuery vs 30M of QTO) and the computationally expensive brute-force approach of QTO that materializes the whole $({\mathcal{V}}\times{\mathcal{V}}\times{\mathcal{R}})$ 3D tensor of scores of all possible triples. Pre-computing such tensors on three datasets takes considerable space and time, e.g., 8 hours for FB15k with heavy sparsification settings to fit onto a 24 GB GPU. Besides, the FB15k dataset is known for years to have leakages (Toutanova & Chen, 2015) from inverse edges that large embedding models memorize and exploit. On a similar note, the graph behind the NELL995 dataset is a collection of disconnected components which is disadvantageous for GNNs.
@@ 5.3 Analysis
Figure 5: Quality analysis on 9 inductive $(e)$ and 3 transductive datasets averaged across all 14 query types. Faithfullness, MRR (left): UltraQuery successfully finds easy answers in larger inference graphs and even outperforms trained GNN-QE baselines. Ranking of easy vs hard answers, ROC AUC (center): zero-shot inference methods slightly lag behind trainable GNN-QE due to assigning higher scores to hard answers. Cardinality Prediction, MAPE (right): UltraQuery is comparable to a much larger trainable baseline QTO. In all cases, UltraQuery LP is significantly inferior to the main model.
@@
Recovering easy answers on any graph. Faithfullness (Sun et al., 2020) is the ability of a CLQA model to return easy query answers, i.e., the answers reachable by edge traversal in the graph without predicting missing edges. While faithfullness is a common problem for many CLQA models, Figure 5 demonstrates that UltraQuery almost perfectly recovers easy answers on any graph size even in the zero-shot inference regime in contrast to the best baseline. Simple score thresholding does not help UltraQuery LP to deal with complex queries as all easy intermediate nodes have high scores above the threshold and the multi-source is more pronounced.
Ranking easy and high answers. A reasonable CLQA model is likely to score easy answers higher than hard ones that require inferring missing links on the fly (Galkin et al., 2022b). Measuring that with ROC AUC, we find (Figure 5) that UltraQuery is about 20% behind the baseline due to less pronounced decision boundaries (overlapping distributions of scores) between easy and hard answers’ scores. However, due to filtering the scores when computing ranking metrics, this fact does not have a direct negative impact on the overall prediction performance.

### 01 Oct 2024 v2

@@ 4.1 Inductive Relation Projection
The vast majority of CLQA methods are inherently transductive and implement relation projections as functions over entity and relation embeddings fixed to a certain KG vocabulary, e.g., with scoring functions from KG completion methods [19, 3, 5], geometric functions [26, 40], or pure neural methods [2, 34]. The only method inductive to new entities [43] learns relation embeddings and uses those as a labeling trick (Section 3) for a GNN that implements the projection operator.
As fixed relation embeddings do not transfer to new KGs with new relations, we adapt Ultra [15], an inductive approach that builds relation representations dynamically using the invariance of relation interactions, as the backbone of the relation projection operator thanks to its good zero-shot performance on simple KG completion tasks across a variety of graphs. Ultra leverages theoretical findings in multi-relational link prediction [6, 20] and learns relation representations from a meta-graph of relation interactions [^2] The meta-graph can be efficiently obtained from any KG.. The meta-graph includes four learnable edge types or meta-relations (head-to-tail, tail-to-head, head-to-head, tail-to-tail) which are independent from KG’s relation vocabulary and therefore transfer across any graph. Practically, given a graph $\mathcal{G}$} and projection query $(h,r,?)$, Ultra employs labeling trick GNNs on two levels. First, it builds a meta-graph ${\mathcal{G}}_{r}$ of relation interactions (a graph of relations where each node is a unique edge type in $\mathcal{G}$}) and applies a labeling trick to initialize the query node $r$. Note that $|{\mathcal{R}}|\ll|{\mathcal{E}}|$, the number of unique relations is much smaller than number of entities in any KG, so processing this graph of relations introduces a rather marginal computational overhead. Running a message passing GNN over ${\mathcal{G}}_{r}$ results in conditional relation representation which are used as initial edge type features in the second, entity-level GNN. There, a starting node $h$ is initialized with a query vector from the obtained relation representations and running another GNN over the entity graph (with a final sigmoid readout) returns a scalar score in $[0,1]$ representing a probability of each node to be a tail of a query $(h,r,?)$.
The only learnable parameters in Ultra are four meta-relations for the graph of relations and GNN weights. The four meta-relations represent structural patterns and can be mined from any multi-relational KG independent of their entity and relation vocabulary. GNN weights are optimized during pre-training. Since the model does not rely on any KG-specific entity or relation vocabulary, a single pre-trained Ultra model can be used as a zero-shot relation projection operator on any KG. Figure 3(a) illustrates the intersection-projection query execution process where each projection step is tackled by the same inductive projection operator with initialization depending on the start anchor node or intermediate variables.
@@ 4.2 Inductive Logical Operations
Learnable logical operators parameterized by neural nets in many CLQA approaches [19, 26, 40, 2] fit a particular embedding space and are not transferable. Instead, we resort to differentiable but non-parametric fuzzy logics [31] that implement logical operators as algebraic operations (t-norms for conjunction and t-conorms for disjunction) in a bounded space $[0,1]$ and are used in several neuro-symbolic CLQA approaches [3, 43, 4, 5, 36]. UltraQuery employs fuzzy logical operators over fuzzy sets ${\bm{x}}\in[0,1]^{|{\mathcal{V}}|}$ as the relation projection operator assigns a scalar in range $[0,1]$ for each entity in a graph. The choice of a fuzzy logic is often a hyperparameter although van Krieken et al. [31] show that the product logic is the most stable. In product logic, given two fuzzy sets ${\bm{x}},{\bm{y}}$, conjunction is element-wise multiplication ${\bm{x}}\odot{\bm{y}}$ and disjunction is ${\bm{x}}+{\bm{y}}-{\bm{x}}\odot{\bm{y}}$. Negation is often implemented as $\mathbf{1}-{\bm{x}}$ where $\mathbf{1}$ is the universe vector of all ones. For second- and later $i$-th hop projections, we obtain initial node states ${\bm{h}}_{v}$ by weighting a query vector ${\bm{r}}_{i}$ with their probability score $x_{v}$ from the fuzzy set of a previous step: ${\bm{h}}_{v}=x_{v}{\bm{r}}_{i}$.
@@ 4.3 Training
Following existing works [25, 43], UltraQuery is trained on complex queries to minimize the binary cross entropy loss
$\hskip-5.0pt\displaystyle{{\mathcal{L}}=-\frac{1}{|{\mathcal{A}}_{q}|}\sum_{a\in{\mathcal{A}}_{q}}\log p(a|q)-\frac{1}{|{\mathcal{V}}\backslash{\mathcal{A}}_{q}|}\sum_{a^{\prime}\in{\mathcal{V}}\backslash{\mathcal{A}}_{q}}\log(1-p(a^{\prime}|q))}$ (1)
@@ 5 Experiments
Our experiments focus on the following research questions: (1) How does a single UltraQuery model perform in the zero-shot inference mode on unseen graphs and queries compared to the baselines? (2) Does UltraQuery retain the quality metrics like faithfullness and identify easy answers reachable by traversal? (3) How does the multi-source propagation issue affect the performance?
@@ 5.1 Setup and Datasets
  • Transductive (3 datasets) where training and inference graphs are the same $(\mathcal{G}_{\textit{train}}=\mathcal{G}_{\textit{inf}})$ and test queries cover the same set of entities and relations: FB15k237, NELL995 and FB15k all from Ren and Leskovec [25] with at most 100 answers per query.
  • Inductive entity $(e)$ (9 datasets) from Galkin et al. [14] where inference graphs extend training graphs $(\mathcal{G}_{\textit{train}}\subset\mathcal{G}_{\textit{inf}})$ being up to 550% larger in the number of entities. The set of relations is fixed in each training graph and does not change at inference making the setup inductive with respect to the entities. Training queries might have more true answers in the extended inference graph.
  • Inductive entity and relation $(e,r)$ (11 datasets): we sampled a novel suite of WikiTopics-QA datasets due to the absence of standard benchmarks evaluating the hardest inductive setup where inference graphs have both new entities and relations $(\mathcal{G}_{\textit{train}}\neq\mathcal{G}_{\textit{inf}})$. The source graphs were adopted from the WikiTopics datasets [16], we follow the BetaE setting when sampling 14 query types with at most 100 answers. More details on the dataset creation procedure are in Appendix A.
Implementation and Training. UltraQuery was trained on one FB15k237 dataset with complex queries for 10,000 steps with batch size of 32 on 4 RTX 3090 GPUs for 2 hours (8 GPU-hours in total). We initialize the model weights with an available checkpoint of Ultra reported in Galkin et al. [15]. Following the standard setup in the literature, we train the model on 10 query types and evaluate on all 14 patterns. We employ product t-norm and t-conorm as non-parametric fuzzy logic operators to implement conjunction $(\wedge)$ and disjunction $(\lor)$, respectively, and use a simple 1x1-x negation. For the ablation study, UltraQuery LP uses the same frozen checkpoint (pre-trained on simple 1p link prediction) with scores thresholding to alleviate the multi-source propagation issue (Section 4.1). More details on all hyperparameters are available in Appendix B.
@@
As common in the literature [25, 27], the answer set of each query is split into easy and hard answers. Easy answers are reachable by graph traversal and do not require inferring missing links whereas hard answers are those that involve at least one edge to be predicted at inference. In the rank-based evaluation, we only consider ranks of hard answers and filter out easy ones and report filtered Mean Reciprocal Rank (MRR) and Hits@10 as main performance metrics.
Other qualitative metrics include: (1) faithfullness [28], i.e., the ability to recover easy answers reachable by graph traversal. Here, we follow the setup in Galkin et al. [14] and measure the performance of training queries on larger inference graphs where the same queries might have new true answers; (2) the ROC AUC score to estimate whether a model ranks easy answers higher than hard answers – we compute ROC AUC over unfiltered scores of easy answers as positive labels and hard answers as negative. (3) Mean Absolute Percentage Error (MAPE) [43] between the number of answers extracted from model’s predictions and the number of ground truth answers (easy and hard combined) to estimate whether CLQA models can predict the cardinality of the answer set.
Baselines. In transductive and inductive $(e)$ datasets, we compare a single UltraQuery model with the best reported models trained end-to-end on each graph (denoted as Best baseline in the experiments): QTO [5] for 3 transductive datasets (FB15k237, FB15k, and NELL995) and GNN-QE [14] for 9 inductive $(e)$ datasets. While a single UltraQuery model has 177k parameters, the baselines are several orders of magnitude larger with a parameters count depending on the number of entities and relations, e.g., a QTO model on FB15k237 has 30M parameters due to having 2000dd entity and relation embeddings, and GNN-QE on a reference FB 175% inductive $(e)$ dataset has 2M parameters. For a newly sampled suite of 11 inductive $(e,r)$ datasets, we compare against the edge-type heuristic baseline introduced in Galkin et al. [14]. The heuristic selects the candidate nodes with the same incoming relation as the last hop of the query. More details on the baselines are reported in Appendix B
5.2 Main Experiment: Zero-shot Query Answering
@@
Model Inductive $(e,r)$ (11 datasets) Inductive $(e)$ (9 datasets) Transductive (3 datasets) Total Average (23 datasets) EPFO avg neg avg EPFO avg neg avg EPFO avg neg avg EPFO avg neg avg MRR H@10 MRR H@10 MRR H@10 MRR H@10 MRR H@10 MRR H@10 MRR H@10 MRR H@10 Best baseline 0.014 0.029 0.004 0.007 0.328 0.469 0.176 0.297 0.468 0.603 0.259 0.409 0.196 0.276 0.105 0.173 UltraQuery 0-shot 0.280 0.380 0.193 0.288 0.312 0.467 0.139 0.262 0.411 0.517 0.240 0.352 0.309 0.432 0.178 0.286 UltraQuery LP 0-shot 0.268 0.409 0.104 0.181 0.277 0.441 0.098 0.191 0.322 0.476 0.150 0.263 0.279 0.430 0.107 0.195 UltraQuery LP no thrs. 0.227 0.331 0.080 0.138 0.246 0.390 0.085 0.167 0.281 0.417 0.127 0.223 0.242 0.367 0.088 0.161
Figure 4: Mitigation of the multi-source message passing issue (Section 4) with UltraQuery: while UltraQuery LP (pre-trained only on 1p link prediction) does reach higher 1p query performance (center right), it underperforms on negation queries (center left). UltraQuery adapts to the multi-source message passing scheme and trades a fraction of 1p query performance for better averaged EPFO, e.g., on the 3i query (right), and negation queries performance. More results are in Appendix C.
@@
Averaged across 23 datasets, UltraQuery outperforms available baselines by relative 50% in terms of MRR and Hits@10 on EPFO and 70% on negation queries (e.g., 0.31 vs 0.20 MRR on EPFO queries and 0.178 vs 0.105 on negation queries). The largest gains are achieved on the hardest inductive $(e,r)$ datasets where the heuristic baseline is not able to cope with the task. On inductive $(e)$ datasets, UltraQuery outperforms the trainable SOTA GNN-QE model on larger inductive inference graphs and performs competitively on smaller inductive versions. On transductive benchmarks, UltraQuery lags behind the SOTA QTO model which is expected and can be attributed to the sheer model size difference (177k of UltraQuery vs 30M of QTO) and the computationally expensive brute-force approach of QTO that materializes the whole $({\mathcal{V}}\times{\mathcal{V}}\times{\mathcal{R}})$ 3D tensor of scores of all possible triples. Pre-computing such tensors on three datasets takes considerable space and time, e.g., 8 hours for FB15k with heavy sparsification settings to fit onto a 24 GB GPU. Still, UltraQuery outperforms a much larger QTO model on the FB15k dataset on both EPFO and negation queries. The graph behind the NELL995 dataset is a collection of disconnected components which is disadvantageous for GNNs.
@@ 5.3 Analysis
Figure 5: Qualitative analysis on 9 inductive $(e)$ and 3 transductive datasets averaged across all 14 query types. Faithfullness, MRR (left): UltraQuery successfully finds easy answers in larger inference graphs and outperforms trained GNN-QE baselines. Ranking of easy vs hard answers, ROC AUC (center): zero-shot inference methods slightly lag behind trainable GNN-QE due to assigning higher scores to hard answers. Cardinality Prediction, MAPE (right): UltraQuery is comparable to a much larger trainable baseline QTO. In all cases, UltraQuery LP is significantly inferior to the main model.
@@
Recovering easy answers on any graph. Faithfullness [28] is the ability of a CLQA model to return easy query answers, i.e., the answers reachable by edge traversal in the graph without predicting missing edges. While faithfullness is a common problem for many CLQA models, Figure 5 demonstrates that UltraQuery almost perfectly recovers easy answers on any graph size even in the zero-shot inference regime in contrast to the best baseline. Simple score thresholding does not help UltraQuery LP to deal with complex queries as all easy intermediate nodes have high scores above the threshold and the multi-source is more pronounced.
Ranking easy and hard answers. A reasonable CLQA model is likely to score easy answers higher than hard ones that require inferring missing links [14]. Measuring that with ROC AUC (Figure 5), UltraQuery is behind the baseline due to less pronounced decision boundaries (overlapping distributions of scores) between easy and hard answers’ scores. Still, due to scores filtering when computing ranking metrics, this fact does not have a direct negative impact on the overall performance.
@@
Figure 6: Average MRR (left) and Hits@10 (right) of 9 inductive $(e)$ and 11 inductive $(e,r)$ CLQA datasets for EPFO and negation queries depending on the number of graphs in the training mix.
Varying the number of graphs in training. Figure 6 and Table 3 report the inductive inference CLQA performance depending on the number of KGs in the training mixture. The original UltraQuery was trained on queries from the FB15k237. In order to maintain the zero-shot inductive inference setup on 11 inductive $(e,r)$ and 9 inductive $(e)$ datasets, we trained new model versions on the rest of BetaE datasets, that is, UltraQuery 2G combines FB15k237 and NELL995 queries (trained for 20k steps), UltraQuery 3G combines FB15k273, NELL995, and FB15k queries (trained for 30k steps). The most noticeable improvement of 2G and 3G versions is the increased MRR and Hits@10 on EPFO queries (9 query types) on 11 inductive $(e,r)$ datasets yielding about 10% gains. On 9 inductive $(e)$ datasets the performance is either on par with the 1G version or a bit lower. Averaged across 20 datasets, the 2G version exhibits the best EPFO performance at the cost of slightly reduced negation query performance.
Table 3: Zero-shot inference results (on 20 inductive datasets) of UltraQuery trained on 1, 2, and 3 datasets, respectively. The biggest gains of the 2G model are in bold.
Model Inductive $(e,r)$ (11 datasets) Inductive $(e)$ (9 datasets) Total Average (20 datasets) EPFO avg neg avg EPFO avg neg avg EPFO avg neg avg MRR H@10 MRR H@10 MRR H@10 MRR H@10 MRR H@10 MRR H@10 UltraQuery 1G 0.280 0.380 0.193 0.288 0.312 0.467 0.139 0.262 0.296 0.423 0.166 0.275 UltraQuery 2G 0.310 0.413 0.187 0.275 0.307 0.463 0.130 0.244 0.308 0.438 0.158 0.260 UltraQuery 3G 0.304 0.402 0.195 0.292 0.292 0.438 0.127 0.239 0.298 0.420 0.161 0.265